### v2

Abstract
With the advancements in quantum computing, transitioning to post-quantum cryptography (PQC) is becoming increasingly critical to maintain the security of modern systems. This paper introduces a formal definition of the cryptographic migration problem and explores its complexities using a suitable directed graph model. Characteristics of the resulting migration graphs are analyzed and trade-offs discussed. By using classical mathematical results from combinatorics, probability theory and combinatorial analysis, we assess the challenges of migrating “random” large cryptographic IT-infrastructures. We show that any sufficiently large migration project that follows our model has an intrinsic complexity, either due to many dependent (comparatively easy) migration steps or due to at least one complicated migration step. This proves that in a suitable sense cryptographic migration is hard in general. Furthermore, we analyze the proposed model with respect to practical applicability and explain the difficulties that emerge when we try to model real-world migration projects.
Keywords:
Cryptographic migration, post-quantum cryptography, hardness results
1 Introduction
Cryptography is fundamental ensuring information security in today’s digital world. Despite advances in cryptographic technologies, the process of migrating existing systems and IT-infrastructures to new standards presents significant challenges [5]. Examples include the transition from IPv4 to IPv6, which took more than 20 years due to the complexity, and technologies such as gateways and NAT became popular migration tools to circumvent the need to actually migrate from IPv4 to IPv6. A cryptographic example is the update of key sizes (RSA keys got bigger and so did the integration problems) or the transition between different hash functions (SHA-1 to the SHA-2 family to the SHA-3 family) with very slow adoption of the SHA-3 family as it is seen as an add-on or fallback rather than a replacement. For a rapid transition, it seems that there needs to be a strong incentive in terms of the benefits of adopting the new technology. One notable example of quicker migration, at least in the broad public, is the transition from the Data Encryption Standard (DES) to the Advanced Encryption Standard (AES), which was driven by the need for stronger security measures due to DES’s vulnerability to brute force attacks [22].
This issue is particularly prominent in the current transition towards PQC, which aims to protect data against future quantum computing threats. The German Federal Office for Information Security (BSI) uses the work-hypothesis for the high-security sector that with significant probability there is a cryptographically relevant quantum computer in the early 2030s [11]. Thus, there are numerous practical post-quantum migrations going on world-wide. The reason is that we are already running-out of time: Sensitive (e. g. classified) data has long confidentiality requirements (e. g. 25 years). In addition migrating the IT-infrastructure correspondingly takes a considerable amount of time. This runs against estimates on how long it will take to build a cryptanalytically relevant quantum-computer, c. f. Mosca’s famous inequality [16]. To understand the structure of cryptographic migration and to potentially improve the practicability of real-world migrations, we will look at the cryptographic migration problem, i. e. the task of migrating a given IT-infrastructure from one cryptographic algorithm to another one, from a mathematical point of view. Despite many common-sense approaches being applied to tackle PQC migration, there remains a significant gap in scientifically grounded methodologies. We address that gap by providing a formal model that lays the foundation for a more rigorous understanding and systematic approach to the complexities of cryptographic migration.
This paper is structured as follows: An overview of related work can be found in Section 2. Our first contribution is to provide a unified and consistent notation by introducing a formal model of cryptographic migration in Section 3. In Section 4 we prove that migration is generally hard in a suitable sense: Namely, we show that any “random”, reasonably large migration project has an inherent complexity. This is our second contribution. Sections 5 and 6 further describe our work in terms of its usage and its potential applications, which shall serve as a foundation for upcoming real-world migrations – our third contribution.
2 Background and Related Work
We will model the cryptographic migration problem as a suitable graph problem. Directed graphs have a long and rich history in computer sciences, see e. g. classical standard references such as [4, 14]. They are used in a wide range of domains due to their ability to model asymmetric relationships and dependencies. For instance, they are used to represent the information exchange in distributed systems [1], the dynamics of the world wide web [28], and communication protocols in network design [26]. To properly place directed graphs in the context of cryptographic migration, we first examine the current state: After the announcements of the first post-quantum public key algorithms for consideration as an international standard by the National Institute of Standards and Technology (NIST) [2], the transition to PQC has started in practice: in the US, there is a large project run by NIST with a large consortium of industry-leading companies that explores the transition to PQC [6]. In Europe, the BSI published a guideline concerning post-quantum migration [10] and also included certain post-quantum algorithms in their recommended suite of cryptographic algorithms [12]. The Dutch government also published a “PQC Migration Handbook” containing guidelines for migrating to PQC [3]. Despite these world-wide activities, we found little research on the structure of migration projects. In fact, many people consider the actual migration of soft- and hardware components somewhat as a best-practice task, and people legitimately ask “Where Is the Research on Cryptographic Transition and Agility?” as David Ott did at RWC2022 (see also [20]). But is migration really a best-practice task? Shouldn’t there be structural reasons why cryptographic migration is so hard to achieve in practice? There are a few informal models of the migration problem (e. g. [21, 18, 27, 9]) but formal treatment is, as far as the authors know, still pending. To objectify this knowledge, we conducted a systematic literature review on this subject [17] without finding a formal model comparable to the one proposed here. This does not completely rule out that there is literature around we did not consider, but we assume that our model can substantially contribute to the formal basics of cryptographic migration.
3 A Formal Model of the Migration Problem
[…]
$v_{10}$$v_{11}$$v_{12}$$v_{9}$$v_{8}$$v_{7}$$v_{6}$$v_{5}$$v_{4}$$v_{2}$$v_{3}$$v_{1}$
Figure 1: An example of a simple migration graph.
We state the important
Definition 1 (Migration Cluster).
Let $G=(V,E)$ be a migration graph and $v\in V$. The migration cluster $c(v)$ of $v$ is the set of all components $w\in V$ such that $w\in\mathop{\text{dep}}\nolimits(v)$ and $v\in\mathop{\text{dep}}\nolimits(w)$.
In the literature, the migration clusters are classically called the strongly connected components of $G$. For sake of exposition we stick to the term migration cluster in the context of cryptographic migration. Intuitively, the elements of $c(v)$ are precisely those components that have to be migrated at the same time as $v$ due to mutual dependencies. As it will turn out, the migration clusters specify how complex single migration steps of a migration project will be. Additionally, the properties of the migration clusters inside a migration graph $G$ will shed light on the complexity of the resulting migration problem which we will describe below. In fact, we have the following important characteristics:
  1. 1.
    The size $s$ of the cluster $s(v):=\#c(v)$.
  2. 2.
    The degree of the cluster $\mathop{\text{cdeg}}\nolimits(v):=\mathop{\text{argmax}}\nolimits_{w\in G|_{c(v)}}\mathop{\text{deg}}\nolimits(w)$.
Clearly, in the case $\mathop{\text{cdeg}}\nolimits(v)=0$ we have $s(v)=1$ since otherwise there would be two components $v_{1}\neq v_{2}\in c(v)$ with $v_{1}\in\mathop{\text{dep}}\nolimits(v_{2})$ and $v_{2}\in\mathop{\text{dep}}\nolimits(v_{1})$. But then $\mathop{\text{dep}}\nolimits(v_{1})=\mathop{\text{target}}\nolimits(v_{1}\to^{*}c(v_{2}))$ would be non-empty and thus $v_{1}\to^{*}c(v_{2})$ would also be non-empty. Therefore, we have $\mathop{\text{deg}}\nolimits_{G|_{c(v_{1})}}(v_{2})>0$, contradicting our assumption. Conversely, if $s(v)=1$ then $\mathop{\text{cdeg}}\nolimits(v)=0$ since $G|_{c(v)}$ consists of a single component only and has no edges. In practice this case occurs in the situation where we have a single component that simply can be migrated without any dependencies on other components. This is the easiest case in a migration process.
In the case $\mathop{\text{cdeg}}\nolimits(v)=1$ the migration cluster $c(v)$ is a cycle. In practice this case often occurs for $s(v)=2$ when there are two components that communicate using the same protocol. A migration of the protocol can only be done successfully, if both components are migrated at the same time. Cycles of larger sizes are uncommon in practice, as we will later see in Section 5. The set $m(v)$ of migratable components around $v$ is given by
$m(v)=\begin{cases}c(v)&\text{, if $c(v)\to\bar{c}(v)=\emptyset$}\\ \emptyset&\text{, otherwise}\\ \end{cases}$
where we write $\bar{c}(v)=V\setminus c(v)$. The set $m(v)$ specifies exactly which components can be migrated in $G$ when we migrate $v\in V$. If $v$ has dependencies outside its migration cluster, i. e. $v$ cannot be migrated, then $m(v)=\emptyset$. If there are no direct dependencies of $v$, then $m(v)=\left\{v\right\}$. Also, in all other cases $m(v)=c(v)$, i. e. the whole migration cluster. If $m(v)\neq\emptyset$, we call $v$ migratable (in $G$).
Now, in a migration project, there might be several subsequent steps to fully migrate all involved components, i. e. solve the migration problem the migration graph poses. We write $G^{(v)}=(V\setminus m(v),E\setminus(V\to m(v)))$ for the graph after a migration attempt of $v$. If $v$ is not migratable, then $G^{(v)}=G$. Generally, $G^{(v)}$ is exactly the graph restricted to the components outside the migratable components around $v$. Using similar to above the notation $\bar{m}(v)=V\setminus m(v)$, we can thus alternatively write $G^{(v)}=G|_{\bar{m}(v)}=(\bar{m}(v),\bar{m}(v)\to\bar{m}(v))$. We now show
Lemma 1.
In any (finite and non-empty) migration graph $G=(V,E)$ there is at least one component $v\in V$ that is migratable.
Proof.
Suppose there is no migratable component, i. e. $m(v)=\emptyset$ for all $v\in V$. By definition, $c(v)\to\bar{c}(v)$ is always non-empty, i. e. there is always a dependency to a component outside the migration cluster of $v$. Call such a dependency $w\in\bar{c}(v)$. Since $w$ is outside the migration cluster of $v$, $w$ cannot depend on elements of $c(v)$ and all dependencies of $w$ reside in the (strict) sub-graph $G|_{\bar{c}(v)}$. Repeating this process, after a finite number of steps, we end-up in a graph which only contains a single migration cluster $c(u)$ and $u\in V$ is migratable in $G$ contradicting our assumption. ∎
We will now define migration strategies of $G$. Formally, a strategy $S=(v_{1},\ldots,v_{\ell})\in V^{\ell}$ for $\ell\in\mathbb{N}$ is a sequence of components that are migrated one by one (of course together with their respective migration clusters). $\ell$ is the length of the strategy. For $1\leq i\leq\ell$ denote by $G_{i}$ the graph after successively migrating $v_{1},\ldots,v_{i}$, i. e. $G_{i}:=G_{i-1}^{(v_{i})}$, where $G_{0}=G$. The strategy $S$ is reasonable if for all $1\leq i\leq\ell$, we have $m(v_{i})\neq\emptyset$ in the graph $G_{i-1}$. That means that the component $v_{i}$ is in fact migratable in the graph $G_{i-1}$ and, therefore $G_{i}\subsetneq G_{i-1}$, i. e. the migration of $v_{i}$ yields a smaller graph. $S$ is successful, if $G_{\ell}$ is the empty graph. By Lemma 1 there is a reasonable, successful strategy for any migration graph $G$.
Lemma 2.
All reasonable, successful strategies $S=(v_{1},\ldots,v_{\ell})$ for a migration graph $G=(V,E)$ have the same length $\ell$.
Proof.
Consider two reasonable, successful strategies $S_{1}=(v_{1}^{1},\ldots,v_{\ell_{1}}^{1})$ and $S_{2}=(v_{1}^{2},\ldots,v_{\ell_{2}}^{2})$ with $\ell_{2}\geq\ell_{1}$. Since $S_{1}$ is successful, $V=\biguplus_{1\leq i\leq\ell_{1}}m(v_{i})$ partitions into the migration clusters given by $S_{1}$. Note that all $m(v_{i})=c(v_{i})$ are non-empty since $S_{1}$ is reasonable. Now, consider the (sub-)strategy $S_{3}=(v_{1}^{2},\ldots,v_{\ell_{1}}^{2})$ of the strategy $S_{2}$. By the partition property, there is for each $1\leq j\leq\ell_{1}$ an $1\leq i\leq\ell_{1}$ such that $v_{j}^{2}\in c(v_{i}^{1})$. Furthermore, for each $1\leq j_{1}<j_{2}\leq\ell_{1}$ and corresponding $1\leq i_{1},i_{2}\leq\ell_{1}$, we have $i_{1}\neq i_{2}$ since otherwise $v_{j_{2}}^{2}\in m(v_{j_{1}}^{2})\neq\emptyset$. The non-emptiness of $m(v_{j_{1}}^{2})$ follows from the assumption that $S_{2}$ is reasonable. But then $\sum_{1\leq j\leq\ell_{1}}s(v_{j}^{2})=\#V$ and the strategy $S_{3}$ is already successful. ∎
Due to Lemma 2 the migration length $\ell(G)$ as a property of the migration graph $G$ is well-defined. It basically expresses how many independent migration steps one has to perform in order to fully migrate a system. Equivalently, we see that the set of components $V$ partitions into the migration clusters of its components, i. e. $V=\bigcup_{v\in V}c(v)$ and for all $v_{1},v_{2}\in V$, we either have $c(v_{1})=c(v_{2})$ or $c(v_{1})\cap c(v_{2})=\emptyset$. In other words the relation $v_{1}\sim_{c}v_{2}:\Leftrightarrow c(v_{1})=c(v_{2})$ is an equivalence relation. Informally, we have
Fact 1.
Migration attempts of parts of an IT-infrastructure cannot increase the difficulty of migrating other, independent parts of the IT-infrastructure.
$c_{9}$$c_{8}$$c_{7}$$c_{6}$$c_{5}$$c_{4}$$c_{2}$$c_{3}$$c_{1}$
Figure 2: The condensation graph $G/c$ for the migration graph from Fig. 1.
[…]
Lemma 3.
For a migration graph $G=(V,E)$ the graph $G/c$ is acyclic.
Proof.
Assume $G/c$ is not acyclic, that is, there are $v_{1},v_{2}\in V$ such that $c(v_{1})\to^{+}c(v_{2})\in(E/c)^{+}$ and $c(v_{2})\to^{+}c(v_{1})\in(E/c)^{+}$ as usual denoting by the superscript $+$ the irreflexive transitive closure of the dependency relation. But then, by definition of $c$, we have $c(v_{1})=c(v_{2})$. ∎
[…]
Figure 3: The migration graph from Fig. 1 with its migration clusters and canonical migration times. In the example with $12$ components, we have $\ell(G)=9$ and $d(G)=5$.
[…]
  1. 1.
    $G$ partitions into the migration clusters of its components.
  2. 2.
    The migration clusters have only acyclic dependencies between each other.
  3. 3.
    The migration clusters have thus a canonical position in the migration process depending on how deep in the graph they are located.
Note that there are also non-canonical migration strategies, where a migration of a component $v\in V$ removed a dependency of another component $w\in V$ and even though there were other migratable components in the first place the migration of $w$ is done next. We will in the following analyze the structure of such graphs for “random migration projects”, i. e. for suitably constructed “random graphs”.
4 Migration is Hard in General
We now show that, expectedly, in any migration graph of certain size there is at least one point in the migration process which is hard to perform. As it turns out there are also relevant trade-offs in any graphs.
4.1 Simple Trade-Offs
First of all, there is a simple trade-off between the number of migration clusters in a graph $G$ and the size of the clusters:
Theorem 4.1.
Assume that for a migration graph $G=(V,E)$ the migration length is bounded by a constant $k$, i. e. $\ell(G)\leq k$. Then, there is a component $v\in V$ with #$\#c(v)\geq n/k$.
Proof.
By definition $\ell(G)=\#(V/c)$ and $n=\sum_{c(u)\in V/c}\#c(u)$. Let $v$ be a component with a maximally large migration cluster, i. e. $v=\mathop{\text{argmax}}\nolimits_{u\in V}\#c(u)$. In general $v$ is not uniquely defined, but for our existential statement, this does not trouble us: Namely, for such $v\in V$, we have $n=\sum_{c(u)\in V/c}\#c(u)\leq\sum_{c(u)\in V/c}\#c(v)=\ell(G)\cdot\#c(v)\leq k\cdot\#c(v)$ and the claim follows. ∎
Informally, we have
Fact 2.
In any sufficiently large migration project, we either have many different migration clusters or at least one large one.
Also locally, there is a similar trade-off between the size of clusters and the number of independent components:
Theorem 4.2.
Consider a component $v$ in a migration graph $G$ with $m\in\mathbb{N}_{>0}$ direct dependencies. Assume there are at most $k$ independent clusters within those $m$ components. Then there is a migration cluster of size $s$ with $s\geq m/k$. Conversely, if each migration cluster has size at most $s$ then we have $k\geq m/s$ migration clusters.
Proof.
Assume each migration cluster has size $s<m/k$. Since there are at most $k$ independent clusters, we have at most $k\cdot s<m$ components. Contradiction. ∎
Thus, we see that we have to expect either many independent components or at least one large migration cluster. Another simple trade-off emerges, when we bound the number of descendants of any component in the graph suitably. Then, we get a large migration length as seen from
Theorem 4.3.
Let ${f\colon\mathbb{R}\rightarrow\mathbb{R}}$ be a smooth function in a real variable with $f\in o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$. Assume for any component $v$ in a migration graph $G=(V,E)$ with $n$ components, we have for each $k\geq 0$ the bound #$\#(v\to^{k}V)\leq f(n)$. Then $\ell(G)\geq n/f(n)$. In other words, the number of elements in the set of components with a common ancestor of distance $k$ has a trade-off with the length of the graph.
Proof.
Without loss of generality we can assume that $G$ has a single source $v$. Otherwise take the set of sources and consider the one with the longest path in $v\to^{*}V$. This path has length $\ell(G)$. For $0\leq k<\ell(G)$ define $W_{k}:=v\to^{k}V$. Since $G$ has only a single source, $\biguplus_{0\leq k<\ell(G)}W_{k}=V$, e. g. $n=\#V=\#\biguplus_{0\leq k<\ell(G)}W_{k}=\sum_{0\leq k<\ell(G)}\#W_{k}$. By assumption #$\#W_{k}\leq f(n)$ for all $0\leq k<\ell(G)$. Thus $n\leq\ell(G)\cdot f(n)$ and the claim follows. ∎
Note that the same trade-off exists for the condensed graph $G/c$ bounding the migration depth $d(G)$ of the graph $G$ from below when assuming for all $k$ bounds #$\#T(k)\leq f(n)$ on the number of migration clusters with canonical migration time $k$ as defined on page 3. This leads to the following informal
Fact 3.
In any sufficiently large migration project, if at all points in time the number of migratable migration clusters is small, then migration will take long. Phrased the other way round: If migration can be done in few successive steps, then at least one of the steps must be difficult.
From classical results in combinatorial analysis we know the following about random decompositions of the migration graph in its clusters.
4.2 Random Migration Clusters
Since every migration graph decomposes disjointly into its migration clusters, we now study the expected properties of these migration clusters. Let $n$ be the number of components. There are $B_{n}$ different possible partitions of $n$ components in disjoint subsets, where
$B_{n}=\sum_{k=0}^{n}{n\choose k}B_{k}=\frac{1}{e}\sum_{k=0}^{\infty}\frac{k^{n}}{k!}$
is the $n$th Bell number. More precisely, there are $n\brace k$ partitions of our components in $k$ distinct non-empty subsets, where $n\brace k$ are called the Stirling numbers of the second kind. Obviously, we have $B_{n}=\sum_{1\leq k\leq n}{n\brace k}$.
4.2.1 Sampling Migration Clusters.
Consider now a migration project with $m$ different sub-projects and suppose each of our $n$ components is assigned to one of the $m$ sub-projects uniformly at random. Let $\Pi_{nm}$ be the random variable defining the resulting partition $\pi$ of our components. We have:
$\mathop{\text{prob}}\nolimits(\Pi_{nm}=\pi)=\frac{m^{\underline{k}}}{m^{n}}$
if #$\#\pi=k$, where $m^{\underline{k}}=m\cdot(m-1)\cdots(m-k+1)$ is the falling factorial function. In other words: if there are $k$ different sub-projects with at least one component, then the probability of creating one such partition is given by the formula above. For the distribution of the number of non-empty different sub-projects we have
$\mathop{\text{prob}}\nolimits(\#\pi=k)={n\brace k}\frac{m^{\underline{k}}}{m^{n}}.$
Note that for arbitrary migration graphs, we do not know the distribution of the different partitions. We have the following
Theorem 4.4 (Proposition 2 in [23]).
Let $M$ be the random variable with values in $\mathbb{N}_{\geq 1}$ and suppose given $M=m$ that $n$ balls labeled by $\mathbb{N}_{\leq n}$ are thrown independently and uniformly at random into $m$ boxes. Then the following two statements are equivalent:
  1. 1.
    $M$ is distributed via $\mathop{\text{prob}}\nolimits(M=m)=m^{n}p_{m}/B_{n}$ for $m\in\mathbb{N}_{\geq 1}$, where $p_{m}$ is a probability distribution on $\mathbb{N}_{\geq 0}$ whose first $n$ factorial moments are identically equal to $1$. One may for example take $p_{m}=1/(em!)$.
  2. 2.
    $\Pi_{nM}$ has uniform distribution over the set of all partitions of $n$ components.
The theorem gives a constructive way of uniformly generating a migration cluster: First sample $m\in\mathbb{N}_{\geq 1}$ following the above distribution, then distribute the $n$ components uniformly in $m$ bins. The resulting partition $\pi$ will be a uniformly selected partition of our $n$ components. We will now study the expected properties of the outcome.
4.2.2 Distribution of Random Migration Clusters.
Studying the distribution of random partitions of sets with $n$ elements is a classical problem. Set into the context of migration problems, we state
Theorem 4.5 (Summary of [19]).
The number of migration clusters $M$ is expectedly $\mathsf{E}(M)\sim n/\mathop{\text{log}}\nolimits n$ and the expected size of each migration cluster is $\mathop{\text{log}}\nolimits n$. Also, almost all partitions have approximately $e\mathop{\text{log}}\nolimits n$ distinct cluster sizes.
Furthermore, the expected maximum size of a migration cluster is, for growing $n$, asymptotic to $er-\mathop{\text{log}}\nolimits\sqrt{r}$, where $r\cdot\mathop{\text{exp}}\nolimits(r)=n$ and $r=\mathop{\text{log}}\nolimits n-(1+o(1))\mathop{\text{log}}\nolimits\mathop{\text{log}}\nolimits n\sim\mathop{\text{log}}\nolimits n$ [24]. This means that we typically expect lots of small clusters and we do not expect exceptionally large ones. Asymptotically, we have
Theorem 4.6 (Chapter 4, Theorem 1.1 in [25]).
The distribution of the number of migration clusters $M$ converges to normal distribution with mean $\mu=n/\mathop{\text{log}}\nolimits n$ and standard deviation $\sigma=\sqrt{n}/\mathop{\text{log}}\nolimits n$ for $n\to\infty$.
Results on the speed of convergence are beyond the scope of this article. We now compute an upper bound on the expected number of migration clusters when looking independently at $s$ randomly selected migration projects.
Lemma 4.
The expected maximal number of migration clusters when looking independently at $s$ randomly selected migration projects is bounded above by
$\frac{n}{\mathop{\text{log}}\nolimits n}+\frac{\sqrt{2n\mathop{\text{log}}\nolimits s}}{\mathop{\text{log}}\nolimits n}.$ (1)
Proof.
Let $M_{1},\ldots,M_{s}$ be the resulting numbers of migration clusters and define $X=\mathop{\text{max}}\nolimits_{1\leq i\leq s}M_{i}-\mu$. Then conveniently $\mathsf{E}(X)=0$.
Since for a fixed real constant $t$ the function $x\mapsto\mathop{\text{exp}}\nolimits(tx)$ is convex, by Jensen’s inequality we have:
$\mathop{\text{exp}}\nolimits(t\mathsf{E}(X))$ $\leq\mathsf{E}(\mathop{\text{exp}}\nolimits(tX))=\mathsf{E}(\max_{1\leq i\leq s}(\mathop{\text{exp}}\nolimits(tM_{i})))$
$\leq\sum_{i=1}^{s}\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$
$=s\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$
Here, the last equality follows from the fact that the moment generating function $\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$ for a normal distribution with mean zero is equal to $\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$ and in particular independent of the index-variable $i$. Rewriting this, $\mathsf{E}(X)\leq\frac{\mathop{\text{log}}\nolimits s}{t}+\frac{t\sigma^{2}}{2}$. This bound is minimal at $\frac{\sqrt{2\mathop{\text{log}}\nolimits s}}{\sigma}$ since its derivative in direction $t$ is $\frac{\sigma^{2}}{2}-\frac{\mathop{\text{log}}\nolimits s}{t^{2}}$. Thus, $\mathsf{E}(X)\leq\sigma\sqrt{2\mathop{\text{log}}\nolimits s}$ and the result follows. ∎
In practice, the Lemma implies that we would have to sample exponentially many migration graphs (in the number of components $n$) in order to expectedly observe graphs with asymptotically more than $n/\mathop{\text{log}}\nolimits n$ migration clusters. Thus, the statement from Theorem 4.5 applies in practice all the time, at least for sufficiently large $n$.
It remains to study the dependencies between the migration clusters, i. e. the structure of the condensation modulo the migration clusters.
4.3 Properties of the Condensation Graph
According to Theorem 4.6, we consider a condensation graph, i. e. the directed acyclic graph resulting from the migration graph on $n$ components modulo its strongly connected components, with expectedly $n^{\prime}=n/\mathop{\text{log}}\nolimits n$ vertices. Since we are considering a single migration project, we can assume that this graph is connected. How can we model this graph?
A first idea would be to use the standard Erdős–Rényi model [8], where the graph is constructed by connecting the components independently at random with a fixed probability $p$. One can ensure that the result is acyclic by defining any linear order on the vertices and include a directed edge only if it obeys the selected ordering. Each such edge is included in the graph with probability $p$, independently from every other edge. For such graphs, we have
Theorem 4.7 (Theorem 4.1 in [13], a variant of Theorem 1 in [7]).
Let $G_{n,p_{n}}$ be a random $n$-vertex graph, where each arc is present with probability $p_{n}=\frac{\mathop{\text{log}}\nolimits n+c_{n}}{n}$. Then, for growing $n$, the probability that $G$ is connected is given by
$\lim_{n\to\infty}\mathop{\text{prob}}\nolimits(G_{n,p_{n}}\text{ is connected})=\begin{cases}0&\text{, if $c_{n}\to-\infty$}\\ e^{-e^{-c}}&\text{, if $c_{n}\to c$ constant}\\ 1&\text{, if $c_{n}\to\infty$}\\ \end{cases}$
Note that we can directly apply this Theorem to our directed acyclic graph situation, since such a graph is connected if and only if the undirected variant of it is connected.
As our goal is to prove a lower bound on the number of edges in the condensation graph (with $n^{\prime}$ vertices!) in our model, we assume now the sparsest case $c_{n^{\prime}}\to c$ constant for which the graph is connected with high probability. Note that the doubly-exponential term in Theorem 4.7 approaches $1$ already for comparatively small $c$ very quickly. For example, already for $c=10$, the value of $e^{-e^{-c}}\geq 0.99995$. But also for any $c>0$, we obtain for the number of edges in the condensation graph expectedly asymptotically
$\frac{\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}}{n^{\prime}}\cdot\frac{n^{\prime}\cdot(n^{\prime}-1)}{2}=\frac{1}{2}((n^{\prime}-1)(\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}))\sim\frac{1}{2}n^{\prime}\mathop{\text{log}}\nolimits n^{\prime}.$
That means that even in the sparsest case, we expect a considerable number of dependencies between the migration clusters.
Also, if we bound the number of components with a common ancestor of distance $k$, i. e. #$\#(v\to^{k}V)$ by a small function $f(n)=o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$, we obtain by Theorem 4.3 for the condensation graph expectedly $d(G)\geq\frac{n}{f(n)\mathop{\text{log}}\nolimits n}$, that is, large chains of dependencies of the migration clusters. Of course one can trade depth $d(G)$ vs. the bound $f(n)$, but then the complexity of at least one migration step will increase.
4.4 Discussion
In the previous sections we analyzed how different graph structures affect the properties of our migration problem. To do so, we looked at random migration graphs, arguing that a random migration project should exhibit random clustering and the remaining directed acyclic condensation should also have a random structure. Thereby, we found several mathematical arguments that underline the fact that migration is hard in general. Specifically, for a random (large) migration graph with $n$ components, we showed that expectedly we have:
  1. 1.
    If for each node the number of dependencies is suitably bounded, the condensation graph has large depth, i. e. many migration clusters that successively depend on each other. Stated oppositely, we have at least one migration step with many dependencies if we assume small depth (Theorems 4.1 and 4.3).
  2. 2.
    Even after observing many migration projects, we do not expect to observe any with considerably more than $n/\mathop{\text{log}}\nolimits n$ migration clusters (Lemma 4).
  3. 3.
    The migration graph decomposes in $n/\mathop{\text{log}}\nolimits n$ migration clusters of size $\mathop{\text{log}}\nolimits n$ with standard deviation $\sqrt{n}/\mathop{\text{log}}\nolimits n$ (Theorem 4.5).
  4. 4.
    The condensation graph is connected and thus not sparse (Theorem 4.7).
We thus showed
Fact 4.
Any sufficiently large migration project has intrinsic complexity due to many dependent (comparatively small) migration clusters. The migration of these clusters either takes many successive steps or includes at least one particularly difficult one.
The question remains how reasonable our assumptions on the randomness of the migration graphs are in practice. The first assumption is that in any practical migration project, partitions (i. e. clustering) of components appears uniformly from the set of all partitions. This clearly does not have to be the case. Another assumption is that the structure of the condensation graph follows the Erdős–Rényi model, i. e. that dependencies appear with a fixed probability independently of each other. This also does not have to be the case. Finally, our third assumption is that the two modeling steps are structurally independent which we are unable to prove.
[…]
Table 1: Numerical evaluation of the expected behavior of migration clusters for migration projects with $n$ components rounded to the nearest integer. We see that one does not expect large mutual dependencies but the expected number of independent clusters is vast and the chain-dependencies are also large when we bound the number of migration clusters with fixed canonical migration time by $f(n)=\mathop{\text{log}}\nolimits n$ or $f(n)=\sqrt{n}/\mathop{\text{log}}\nolimits n$, respectively.
We conclude that even though the mathematical model does not represent the exact structure of every migration project, it gives a first sound basis for understanding the structure of migration projects in general. We summarize these results with a numerical evaluation of the relevant estimates in Table 1.
5 Real-World Dependencies
We provided the theoretical foundation for working with migration graphs in general. Now we take a first step towards its application in the real world. One has to consider that a migration graph’s structure will vary significantly depending on the concrete IT-infrastructure. But also the abstraction layer and the type of migration lead to many design choices of the model. The resulting graphs will show parallels to those that already exist in several disciplines of computer science. Those comprise flowcharts, call graphs and finite-state machines in programming, network diagrams, or software dependency graphs, just to name a few. Additionally, current efforts in the creation of Software Bills of Materials (SBOMs) or Cryptography Bills of Materials (CBOMs) and their respective tooling add to that collection of suitable graphs.
5.1 Post-Quantum Migration
We expect that post-quantum migration is the most significant type of cryptographic migration of IT-infrastructures in the near future [17]. Its specific properties introduce aspects that were less prevalent in the classical cryptography world. One example are hybrid cryptographic schemes. Depending on the specific model, a hybrid solution may just be seen as a single cryptographic primitive or cipher suite. Yet, by modeling detailed software dependencies or taking dependencies on space/throughput requirements into account, this might have a significant effect.
Post-quantum migration also emphasizes the importance of cryptographic agility (see for example [20]). While many communication protocols used to be optimized for a limited set of somewhat similar algorithms, the introduction of crypto-agile mechanisms that support various kinds of cryptographic primitives and, sometimes arbitrary, combinations of these, will lead to an increase in complexity. In a graph which only models systems and their protocol-related communication relations in between, one may only treat this implicitly. Yet again, graphs that take other dependencies into account become harder to model depending on the level of detail but, on the other hand, sometimes better reflect the complicated circumstances of the real world.
Also, regulatory requirements of post-quantum solutions and different views on current and future algorithm recommendations may become way more complicated, especially in a multinational context.
All in all, post-quantum migration is the scenario that motivated our formal migration model. At the same time, it implicates several special properties which make modeling even more complex as we demonstrate in the following section.
5.2 Real-World Migration Graphs
Today’s large scale IT-infrastructures are both complex and highly diverse and examining models to create a suitable graph from a real-world IT-infrastructure is an open research topic.
The aggregation of the necessary data that is used to derive the graph is highly individual and specific to the given migration goal. We state a few concrete examples of such modeling attempts:
Network migration:
When talking about network migration it seems intuitive to think of the components in $V$ to represent computers, servers and similar devices. Much thought has to be taken into the dependencies depicted by the edges $E$. These could, e. g. model communication paths for specific (cryptographic) protocols. For the same network all connections between machines may be interesting regardless of the specific protocol, illustrating all the relations between machines in the network. Depending on the objective one may study several graphs representing different kinds of communication relations each – for example individual graphs for protocols like TLS/HTTPS, SSH, IPsec/IKEv2).
System migration:
Different graphs may be modeling a specific system, e. g. $V$ being a set of components within a firmware or operating system. It is unclear at which granularity such modeling may be done. We expect that relations may be handled similarly to SBOMs, showing dependencies between software, libraries, and so on. This may even go into the details of a specific software project showing relations between different modules, library calls, etc.
Full-scale migration:
For a system administrator, a tool focusing on the high-level network view may be enough, while programmers may profit from call-graph-like models. A full view on all abstraction levels may be provided by a combination of different graphs, e. g. a component $v$ being linked to another graph depicting the same component in a different context. This leads us to having differing graph models of the same infrastructure, just depending on our perspective.
Note that when we consider classical, naturally grown enterprise networks, the graphs will probably be both: large and complex. Also they will have many interdependent systems. It is beyond our current model to capture time dependent changes in the migration graph. But even when we look at a snap-shot of a migration graph, the design choices for modeling real-world IT-infrastructure graphs are vast and we discovered a variety of possible ways to model such IT-infrastructures.
A reasonable assumption might be that the best strategy is to model only what is vital for the given migration and leave out all other information. Examining the exact choices, however, deserves its own research track.
5.3 Real-World Patterns
$v$
$v_{1}$$v_{2}$$v_{3}$$w$
$v$$w_{2}$$w_{1}$$w_{3}$
$v$$w$
$v_{1}$$v_{2}$$v_{3}$
Figure 4: Dependency sub-graphs of real-world migration attempts: 4(a) Single component with no dependencies. 4(b) Many components depend on a single component. 4(c) One component depends on many components. 4(d) Mutual dependencies between two components. 4(e) Higher-order cycles.
When analyzing the occurrence of graph patterns in the real world, we might begin with the kinds of typical dependency patterns we observe in practice. To do so, we depicted different base-patterns in Fig. 4, from which any directed graph can be constructed. In the following, we will express real-world observations related to these patterns and discuss their semantics and meaning for our introduced theory. Also, we briefly reflect on their impact on post-quantum migration.
API-compatible change of algorithms:
When migrating an isolated (generic) technological component $v$, a simple change in the code or even an update of the relevant software migrates the component successfully. This corresponds to case Fig. 4(a). In the context of post-quantum migration this could be a simple shift from one cryptographic algorithm to an API-compatible post-quantum algorithm, e. g. the change from ECDSA to CRYSTALS-Dilithium.
Operating System dependencies:
Certain components $v_{1},v_{2},v_{3},\dots$, for example software or libraries, can only be migrated when the underlying execution platform $w$ supports the newly migrated component. This often occurs when one version of an application is updated but the new version only runs on a newer version of the operating system. This corresponds to case Fig. 4(b). In the context of post-quantum migration this could be the replacement of a necessary cryptographic hardware resource running a certain operating system with a PQC-ready version of it.
Regulatory requirements:
When regulatory requirements $w_{1},w_{2},w_{3},\dots$ impose restrictions on the migratability of certain technical components $v$, only a change in the regulatory requirements enables us to actually migrate. Changing regulatory requirements, however, often depends on many further, often non-technical, components, such as laws, standardization, technological standards in general, etc. This corresponds to case Fig. 4(c) and directly applies to post-quantum migration as well.
Communication protocols:
When migrating communication protocols from one version to another, it is vital that all communicating parties are migrated at the same time. This is in particular a problem when there are many parties involved and only a part of the own IT-infrastructure may be updated conveniently, e. g. by software configuration management tools with external dependencies. This problem often leads to backwards compatibility as a feature, but likely also leads to the possibility of downgrading attacks. This corresponds to case Fig. 4(d). In the context of post-quantum migration this could be the replacement of TLS by a post-quantum-ready version of it.
Dependency on space/throughput requirements:
Occasionally, there are requirements on the amount of memory, the capabilities of computing platforms, or network throughput restrictions. This corresponds to cases Figs. 4(d) and 4(e). In the context of post-quantum migration this applies when for example the key sizes of post-quantum algorithms are much larger than those of classical algorithms.
Dependencies of software-packages within distributions:
When considering software-packages and their dependencies, the graph formed this way is a directed acyclic graph with no mutual dependencies. This graph represents combinations of cases illustrated in Figs. 4(a)4(b), and 4(c). The migration of such a case can be done corresponding to the canonical topological order of this graph. In the context of post-quantum migration this could be the replacement of cryptographic libraries with post-quantum-ready versions.
Cyclically depending assets:
The draft of ETSI’s “Repeatable Framework for Quantum-safe Migrations” [9] explains that if for example three cryptographic assets are in a dependency relationship with each other, functionality issues may occur when a single asset is migrated. The functionality issues might cascade and affect all assets in the circle, even the migrated one. This corresponds to case Fig. 4(e), but the problem is mitigated in our model as we assume that such circular clusters of assets can only be migrated when all assets are migrated at the same time. In the context of post-quantum migration this might happen for example when we observe circularly cross-signed certificates. Such circular cases are rare in practice, because they create dependencies that are avoided in software architecture and engineering. Also, implementers can address those kinds of dependencies by introducing adapter nodes in one direction which allow to break the dependency cycle.
We encourage the inclined reader to find more naturally occurring real-world examples. Using these base-patterns, a wide range of possible security scenarios can not only be modeled, but also described and simplified, by breaking them down into smaller, well-understood structures. Thus, once identified, the above patterns might serve as fundamental steps when deriving migration strategies in practical settings.
6 Practical Implications
As stated in Fact 4 and further outlined in Section 5, big migration projects are intrinsically complex. These projects require many successive steps and in-depth strategic planning. Migration graphs can help in understanding this complexity involved and in considering the time needed to migrate, but also specify the number of components and the various dependencies. As we have seen, these dependencies can exist between single components, between interdependent clusters of components, and between external factors such as regulatory requirements, legal issues, or third-party components. The complexity of migrations requires expert knowledge and sufficient dedicated resources, resulting in significant costs. Economic implications and the risks of both, migrating and not migrating to PQC, must be evaluated. It is also crucial for stakeholders, policy makers, and non-technical decision makers to grasp this complexity and to develop appropriate organizational measures to address it.
Ott et al. define the term cryptographic agility as “the ability to make a transition from one cryptographic algorithm or implementation to another” [20]. Such an approach is in line with making security solutions sustainable by allowing to adapt to changing threats and technologies. We have shown that it is necessary to approach a migration process step by step and reevaluate the changed system once the first migrations have been performed. The model we propose can serve as a starting point to identify the highest priority parts of the system, e. g., by identifying the parts which are at the beginning of a migration chain, or which have the highest impact. We believe that our approach works well in an iterative migration approach and we will investigate this further in future work.
Standardizing migration projects through guidelines, rules, and best practices can optimize the process, provide practical examples for guidance, and reduce cost and effort. First steps in these directions are approaches like CARAF, a risk management approach for post-quantum migration [15]. CARAF defines a threat assessment and a risk assessment stage in which migration graphs as we define them can be helpful. During threat assessment, migration clusters identified using our approach can be used to identify vulnerable subsystems. If one of the systems in the cluster is vulnerable, so is the cluster in its entirety. Likewise, during risk assessment, information about which systems need to be migrated together is helpful to establish cost and effort estimates. Therefore, we believe our approach supports existing methods in the area [17].
Currently, our approach is not supported by tools. We believe that the development and refinement of open-source tools and frameworks, along with ensuring easy access to these resources, can significantly lower the barriers for migration, especially for smaller organizations. These tools can streamline migration processes and offer cost-effective solutions for managing complex projects. We envision tooling to support the creation and maintenance of migration graphs as well as their analysis. Similar tools can be found in related areas such as network analysis and visualization, SBOMs and CBOMs, and graph-based databases. The formalization approach may facilitate the use and adaption of such tools by providing a structured understanding of the underlying complexities. Our future work will also address costs and effort associated with migrations and how these can be captured in the graphs.
In summary, managing the intrinsic complexity of large migration projects involves understanding the dependencies and strategically planning of the different migration steps, while keeping the economic implications in mind. Supporting companies with standardized guidelines and open-source tools is crucial, as is fostering an agile approach to sustainable security solutions. By addressing these key areas, organizations can overcome migration challenges and maintain robust security within their IT-infrastructures.
7 Conclusion
In this work, we used classical mathematical results from graph theory, probability theory, and combinatorial analysis to assess the difficulty of a very practical problem: Cryptographic migration of large IT-infrastructures. We formulated a formal model of cryptographic migratability in terms of directed graphs and showed several expected properties of these graphs. This approach not only elevates the discussion to a scientific level but also establishes a foundation for further research on more efficient migration strategies. From these formal results, we derived four “facts” that informally state that migration is difficult in general. In summary, we showed:
  1. 1.
    Migration attempts of independent parts of the IT-infrastructure do not interfere with each other.
  2. 2.
    In sufficiently large migration projects we either have many different migration clusters or at least one large one.
  3. 3.
    In sufficiently large migration projects, if at all points in time the number of migratable migration clusters is small, then migration will take long.
  4. 4.
    Such migration projects always have an intrinsic complexity due to many dependent (comparatively small) migration clusters which (successively) depend on each other.
Still, there is a lot of work to be done, when it comes to practical migration projects: It is not clear how to obtain useful migration graphs in practice, such that they are neither trivial nor too vast. Analyzing how to model real-world IT-infrastructures or relevant subsets will clearly be a challenging task. Also, there is clearly room for improvements: As migration problems typically come with some kind of measurable cost per migration step, a refinement of our model would result in migration graphs where optimal migration strategies are not as easy to obtain as here. One further important open question is also the development of migration graphs and the changes in such graphs over time.
Our work sets the stage for future research aimed at refining these models, integrating them with real-world data, and ultimately creating more adaptive, user-friendly tools that support real-world migration. The paper therefore aims to scientifically substantiate migration approaches. By emphasizing the relevance of this formal approach, we hope to support the development of more robust migration strategies and (open-source) tooling that can keep pace with the evolving landscape of cryptographic standards.
Future work should apply our model to real-world migration projects. The hope is to get a better insight of the complexity of migration projects not just in general, as discussed here, but for specific practically occurring cases. This might help to better understand and categorize real-world migration steps and processes, specifically in the context of PQC migration.
Acknowledgments
The authors thank Tobias Bauer for his valuable comments during the genesis of this article.