### v3

# On the Formalization of Cryptographic Migration

Abstract
With the advancement of quantum computing, the transition to post-quantum cryptography (PQC) is becoming increasingly critical to maintain the security of modern dependable infrastructural systems. This paper presents a novel approach to gain insight into the structure of cryptographic migration problems, using a semi-formal model to capture the inherent dependencies and complexities of such transitions. Using classical mathematical results from combinatorics, probability theory, and combinatorial analysis, we assess the challenges of migrating large cryptographic IT-infrastructures and prove that – in a suitable sense – cryptographic migration has a certain expected complexity. Furthermore, we analyze the proposed model in terms of real-world patterns as well as its practical applicability, and discuss difficulties that arise when trying to model real-world migration projects. This work sets the stage for future advances in both the theoretical understanding and practical implementation of cryptographic migration strategies in the post-quantum era.
Index Terms:
Cryptographic migration, post-quantum cryptography, semi-formal model
I Introduction
Cryptography is fundamental to ensuring the information security in today’s digital world and is thus one important cornerstone of many modern dependable systems. Despite advances in cryptographic technologies, the process of migrating existing systems and IT-infrastructures to new standards presents significant challenges [3]. One example is the transition from IPv4 to IPv6, which took more than 20 years and promoted technologies such as gateways and NAT to circumvent the actual migration instead. A cryptographic example is the update of key sizes (RSA keys got bigger and so did the integration problems) or the transition between different hash functions (SHA-1 to the SHA-2 family to the SHA-3 family) with slow adoption of the SHA-3 family as SHA-2 seems still sufficient for most tasks. It seems that a rapid transition requires a strong incentive in terms of the benefits of adopting the new technology. One notable example of quicker migration, at least in the broad public, is the transition from the Data Encryption Standard (DES) to the Advanced Encryption Standard (AES), which was driven by the need for stronger security measures due to DES’s vulnerability to brute force attacks [20].
The issue of migration is particularly prominent in the current transition towards PQC, which aims to protect data against future quantum computing threats. The German Federal Office for Information Security (BSI) uses the work-hypothesis for the high-security sector that there likely exists a cryptographically relevant quantum computer in the early 2030s [8]. Thus, there are numerous practical post-quantum migrations going on world-wide. The reason is that we are already running-out of time: Sensitive (e. g. classified) data often has long confidentiality requirements (e. g. 25 years). In addition, migrating the IT-infrastructure correspondingly takes a considerable amount of time [24]. This runs against estimates on how long it will take to build a crypto analytically relevant quantum-computer, c. f. Mosca’s famous inequality [13].
To be able to conduct such migrations as efficiently as possible we aim to model the PQC migration problem mathematically and look for potential optimal strategies. As two relevant parts, i.e. the modeling of IT networks and dependencies, are often depicted as graphs [11], we also examine the migration problem in the context of a graph-based model. By this systematic analysis, we hope to better understand the structural problems and find strategies necessary for building tools to assist with conducting the actual migrations. Our findings show that usually the systems’ dependencies of a to-be-migrated infrastructure necessitate a compatible migration order, otherwise functionality is lost. Within this order there are always sets of systems which need to be migrated together – so called migration clusters. As soon as the systems under evaluation exhibit non-trivial complexity, we have many or large clusters, and these clusters are expected to be interdependent, making migration difficult. To facilitate the process and still allow for rapid migration, we see a need for tools that build on our results to find and visualise the optimal migration graphs.
After an overview of related work in Section II, we present our first contribution in Section III. It consists of a formalization approach for cryptographic migration, a discussion of the architectural structure of the resulting model, and explores typical patterns that occur in practice. In Section IV we prove that any (reasonably large) migration project has an expected intrinsic complexity as our second contribution. Section V further describes our model in terms of its usage and its potential applications which shall serve as a foundation for tools which assist the upcoming real-world migrations – our third contribution.
II Background and Related Work
After the announcements of the first post-quantum public key algorithms for consideration as an international standard by the National Institute of Standards and Technology (NIST) [1], the transition to PQC has started in practice: in the US, there is a project run by NIST with a large consortium of industry-leading companies that explores the transition to PQC [3]. In Europe, the German BSI published a guideline concerning post-quantum migration [7] and also included certain post-quantum algorithms in their recommended suite of cryptographic algorithms [9]. The Dutch government also published a “PQC Migration Handbook” containing guidelines for migrating to PQC [2].
Despite these world-wide activities, we found little research on both, the general structure of migration projects and how to suitably model cryptographic migration. In fact, many people consider migrating of soft- and hardware components as a best-practice task. So, people legitimately ask “Where Is the Research on Cryptographic Transition and Agility?” as David Ott did at RWC2022 (see also [18]). Shouldn’t there be structural reasons why cryptographic migration is so hard to achieve in practice? We will show in Section IV that there is!
Recent efforts have focused on consolidating the state-of-the-art in cryptographic migration [14] and systematically organizing knowledge on cryptographic agility [16]. Furthermore, frameworks arise that aim to support the migration process as in [25], and the development of tools to assist in the migration process [11]. The latter follows – as we do – a graph-based approach, providing a solid starting point for PQC migration by introducing a practical framework focused on cryptographic and data inventories and dependency analysis within enterprise systems. However, the analysis in [11] primarily addresses operational aspects and could benefit from a more formal approach. Building on this groundwork, our efforts enhance the theoretical rigor by formalizing migration challenges through mathematical models. Generalizing the approach might even address broader migration scenarios beyond PQC-specific contexts.
Also to be mentioned are few informal models of the migration problem (e. g. [19, 15, 23, 6]). However, formal treatment is, as far as the authors know, still pending. We believe that our migration model together with its analysis can substantially contribute to the formal basics of cryptographic migration.
III A Model for the Cryptographic Migration Problem
Any IT-infrastructure we migrate consists of a number of migration components or simply components which need to be migrated. Such migration components can be software libraries, whole software packages, full operating systems, network protocols, etc. just to name a few. For a more thorough study of which components we might encounter in practice see Section III-B. The migration problem is now to migrate all migration components. However, not every component can be migrated at every point in time, because there might be dependencies between the components. For example, a given software package can only be migrated to post-quantum if its used cryptographic libraries can. Another example is the migration of a communicating software component, which can only be migrated if the corresponding counterparts of the communication are migrated at the same time. In this case the migration can be done (if there are no other dependencies) if all communicating parties of this single migration component are migrated at the same time. We call this as stated in the introduction in Section I a migration cluster. Components without dependencies outside its migration cluster are called migratable components. The migration process looks now as follows:
  1. 1.
    In every migration step we select, following a suitable migration strategy, a migratable component.
  2. 2.
    We then compute its migration cluster, i.e. possible interdependent components that need to be migrated at the same point in time.
  3. 3.
    Then, the actual migration of the migration cluster is performed and the successfully migrated migration cluster is removed from our model (together with all in-going dependencies).
  4. 4.
    This process is repeated until there is no migration component left and we have then solved the migration problem for this specific IT-infrastructure.
The whole model is now captured in a directed graph whose vertices are the migration components and the edges are the dependencies between components, the migration graph of the migration problem, see Appendix -A for the formal details.
Note that a migration graph’s structure will vary significantly depending on the concrete IT-infrastructure and the level of model we have in mind, c.f. Section III-A. But also the abstraction layer and the type of migration lead to many design choices of the model. The resulting graphs will show parallels to those that already exist in several disciplines of computer science. Those comprise flowcharts, call graphs, and finite-state machines in programming, network diagrams, or software dependency graphs, just to name a few. Additionally, current efforts in the creation of Software Bills of Materials (SBOMs) or Cryptography Bills of Materials (CBOMs) and their respective tooling add to that collection of suitable graphs.
We expect that post-quantum migration is the most significant type of cryptographic migration of IT-infrastructures in the near future [14]. Its specific properties introduce aspects that were less prevalent in the classical cryptography world. One example are hybrid cryptographic schemes. Depending on the specific model, a hybrid solution may just be seen as a single cryptographic primitive or cipher suite. Yet, modeling detailed software dependencies or taking dependencies on space/throughput requirements into account might have a significant effect.
Post-quantum migration also emphasizes the importance of cryptographic agility (see for example [18]). Many communication protocols used to be optimized for a limited set of somewhat similar algorithms. The introduction of crypto-agile mechanisms that support various kinds of cryptographic primitives and, sometimes arbitrary, combinations of these, will lead to an increase in complexity. In a graph which only models systems and their protocol-related communication relations in between, one may only treat this implicitly. Yet again, while graphs that take other dependencies into account usually become harder to model, they sometimes better reflect the complicated circumstances of the real world.
Also, regulatory requirements of post-quantum solutions and different views on current and future algorithm recommendations may become significantly more complicated, especially in a multinational context.
All in all, post-quantum migration is the scenario that motivated our formal migration model. At the same time, it implicates several special properties which make modeling even more complex as we demonstrate in the following section.
III-A Real-World Migration Graphs
Today’s large scale IT-infrastructures are both complex and highly diverse, and examining models to create a suitable graph from a real-world IT-infrastructure is an open research topic. The aggregation of the necessary data that is used to derive the graph is highly individual and specific to the given migration goal. We state a few concrete examples of such modeling attempts:
Network migration:
When talking about network migration it seems intuitive to think of the components to represent computers, servers, and similar devices. Much thought has to be taken into the dependencies. These could, e. g. model communication paths for specific (cryptographic) protocols. For the same network, all connections between machines may be interesting regardless of the specific protocol, illustrating all the relations between machines in the network. Depending on the objective one may study several graphs representing different kinds of communication relations each – for example individual graphs for protocols like TLS/HTTPS, SSH, or IPsec/IKEv2.
System migration:
Different graphs may be modeling a specific system, e. g. the migration components being a set of software-artifacts within a firmware or operating system. It is unclear at which granularity such modeling may be done. We expect that relations may be handled similarly to SBOMs, showing dependencies between software, libraries, and so on. This may even go into the details of a specific software project showing relations between different modules, library calls, etc.
Full-scale migration:
For a system administrator, a tool focusing on the high-level network view may be enough, while programmers may profit from call-graph-like models. A full view on all abstraction levels may be provided by a combination of different graphs, e. g. a component being linked to another graph depicting the same component in a different context. This leads us to having differing graph models of the same infrastructure, just depending on our perspective.
Note that when we consider classical, naturally grown enterprise networks, the graphs will probably be both: large and complex. Also, they will have many interdependent systems. It is beyond our current model to capture time dependent changes in the migration graph. But even when we look at a snap-shot of a migration graph, the design choices for modeling real-world IT-infrastructure graphs are vast and we discovered a variety of possible ways to model such IT-infrastructures.
A reasonable assumption might be that the best strategy is to model only what is vital for the given migration and leave out all other information. Examining the exact choices, however, deserves its own research track.
III-B Real-World Patterns
$v$
$v_{1}$$v_{2}$$v_{3}$$w$
$v$$w_{2}$$w_{1}$$w_{3}$
$v$$w$
$v_{1}$$v_{2}$$v_{3}$
Figure 1: Dependency sub-graphs of real-world migration attempts: 1(a) Single component with no dependencies. 1(b) Many components depend on a single component. 1(c) One component depends on many components. 1(d) Mutual dependencies between two components. 1(e) Higher-order cycles.
When analyzing the occurrence of graph patterns in the real world, we might begin with the kinds of typical dependency patterns we observe in practice. To do so, we depicted different base-patterns in Fig. 1, from which any directed graph can be constructed. In the following, we will express real-world observations related to these patterns and discuss their semantics and meaning for our introduced theory. Also, we briefly reflect on their impact on post-quantum migration.
API-compatible change of algorithms:
When migrating an isolated (generic) technological component $v$, a simple change in the code or even an update of the relevant software migrates the component successfully. This corresponds to case Fig. 1(a). In the context of post-quantum migration this could be a simple shift from one cryptographic algorithm to an API-compatible post-quantum algorithm, e. g. the change from ECDSA to CRYSTALS-Dilithium.
Operating System dependencies:
Certain components $v_{1}$, $v_{2}$, $v_{3}$, $\dots$, for example software or libraries, can only be migrated when the underlying execution platform $w$ supports the newly migrated component. This often occurs when one version of an application is updated but the new version only runs on a newer version of the operating system. This corresponds to case Fig. 1(b). In the context of post-quantum migration, this could be the replacement of a necessary cryptographic hardware resource running a certain operating system with a PQC-ready version of it.
Regulatory requirements:
When regulatory requirements $w_{1}$, $w_{2}$, $w_{3}$, $\dots$ impose restrictions on the migratability of certain technical components $v$, only a change in the regulatory requirements enables us to actually migrate. Changing regulatory requirements, however, often depends on many further, often non-technical, components, such as laws, standardization, technological standards in general, etc. This corresponds to case Fig. 1(c) and directly applies to post-quantum migration as well.
Communication protocols:
When migrating communication protocols from one version to another, it is vital that all communicating parties in the migration cluster are migrated at the same time. This is in particular a problem when there are many parties involved and only one part of the own IT-infrastructure may be updated conveniently, e. g. by software configuration management tools with external dependencies. This problem often leads to backwards compatibility as a feature, but likely also leads to the possibility of downgrading attacks. This corresponds to case Fig. 1(d). In the context of post-quantum migration this could be the replacement of TLS by a post-quantum-ready version of it.
Dependency on space/throughput requirements:
Occasionally, there are requirements on the amount of memory, the capabilities of computing platforms, or network throughput restrictions. This corresponds to cases Figs. 1(d) and 1(e). In the context of post-quantum migration this applies when for example the key sizes of post-quantum algorithms are much larger than those of classical algorithms.
Dependencies of software-packages within distributions:
When considering software-packages and their dependencies, the graph formed this way is a directed acyclic graph with no mutual dependencies. This graph represents combinations of cases illustrated in Figs. 1(a)1(b), and 1(c). The migration of such a case can be done corresponding to the canonical topological order of this graph. In the context of post-quantum migration this could be the replacement of cryptographic libraries with post-quantum-ready versions.
Cyclically depending assets:
The draft of ETSI’s “Repeatable Framework for Quantum-safe Migrations” [6] explains that if for example three cryptographic assets are in a dependency relationship with each other, functionality issues may occur when a single asset is migrated. The functionality issues might cascade and affect all assets in the circle, even the migrated one. This corresponds to case Fig. 1(e) but the problem is mitigated in our model as we assume that such circular clusters of assets can only be migrated when all assets are migrated at the same time. In the context of post-quantum migration this might happen for example when we observe circularly cross-signed certificates. Such circular cases are rare in practice, because they create dependencies that are avoided in software architecture and engineering. Also, developers can address those kinds of dependencies by introducing adapter nodes in one direction which allow to break the dependency cycle.
We encourage the inclined reader to find more naturally occurring real-world examples. Using these base-patterns, a wide range of possible security scenarios can not only be modeled, but also described and simplified, by breaking them down into smaller, well-understood structures. Thus, once identified, the above patterns might serve as fundamental steps when deriving migration strategies in practical settings.
IV Assessing the Difficulty of the Migration Problem
We now show that, expectedly, any sufficiently large migration project has an intrinsic complexity. Specifically we show:
Theorem 1.
Expectedly we have for migration projects with $n$ migration components within our model:
  1. 1.
    The number of migration steps $M$ is expectedly $\mathsf{E}(M)\sim n/\mathop{\text{log}}\nolimits n$ with standard deviation $\mathsf{SD}(M)\sim\sqrt{n}/\mathop{\text{log}}\nolimits n$ and the expected size of each migration cluster is $\mathop{\text{log}}\nolimits n$.
  2. 2.
    The expected maximal number of migration clusters when looking independently at $s$ randomly selected migration projects is bounded above by
    $\frac{n}{\mathop{\text{log}}\nolimits n}+\frac{\sqrt{2n\mathop{\text{log}}\nolimits s}}{\mathop{\text{log}}\nolimits n}.$
  3. 3.
    If the set of migratable components is at each step bound by a smooth function ${f\colon\mathbb{R}\rightarrow\mathbb{R}}$ in a real variable with $f\in o(n)$, then we expect $n/f(n)$ non-independent migration steps that successively depend on each other.
Proof.
Since every migration project decomposes disjointly into its migration clusters, we now study the expected properties of these migration clusters. Let $n$ be the number of components. There are $B_{n}$ different possible partitions of $n$ components in disjoint subsets, where $B_{n}$ is the $n$th Bell number. Denote by $M$ the number of disjoint subsets of a randomly chosen partition.
  1. 1.
    The statement on the expectation and the standard deviation of $M$ is a summary of [17] and Chapter 4, Theorem 1.1 in [22]. Furthermore, the expected maximum size of a migration cluster is, for growing $n$, asymptotic to $er-\mathop{\text{log}}\nolimits\sqrt{r}$, where $r\cdot\mathop{\text{exp}}\nolimits(r)=n$ and $r=\mathop{\text{log}}\nolimits n-(1+o(1))\mathop{\text{log}}\nolimits\mathop{\text{log}}\nolimits n\sim\mathop{\text{log}}\nolimits n$ [21]. Results on the speed of convergence are beyond the scope of this article.
  2. 2.
    Let $M_{1},\ldots,M_{s}$ be the resulting numbers of migration clusters and define $X=\mathop{\text{max}}\nolimits_{1\leq i\leq s}M_{i}-\mu$. Then conveniently $\mathsf{E}(X)=0$.
    Since for a fixed real constant $t$ the function $x\mapsto\mathop{\text{exp}}\nolimits(tx)$ is convex, by Jensen’s inequality we have:
    $\mathop{\text{exp}}\nolimits(t\mathsf{E}(X))$ $\leq\mathsf{E}(\mathop{\text{exp}}\nolimits(tX))=\mathsf{E}(\max_{1\leq i\leq s}(\mathop{\text{exp}}\nolimits(tM_{i})))$
    $\leq\sum_{i=1}^{s}\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$
    $=s\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$
    Here, the last equality follows from the fact that the moment generating function $\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$ for a normal distribution with mean zero is equal to $\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$ and in particular independent of the index-variable $i$. Rewriting this, $\mathsf{E}(X)\leq\frac{\mathop{\text{log}}\nolimits s}{t}+\frac{t\sigma^{2}}{2}$. This bound is minimal at $\frac{\sqrt{2\mathop{\text{log}}\nolimits s}}{\sigma}$ since its derivative in direction $t$ is $\frac{\sigma^{2}}{2}-\frac{\mathop{\text{log}}\nolimits s}{t^{2}}$. Thus, $\mathsf{E}(X)\leq\sigma\sqrt{2\mathop{\text{log}}\nolimits s}$ and the result follows.
  3. 3.
    We study the dependencies between the migration clusters. According to the first statement of this Theorem, we consider the directed acyclic graph resulting from the migration graph on $n$ components modulo its strongly connected components with expectedly $n^{\prime}=n/\mathop{\text{log}}\nolimits n$ vertices. Since we are considering a single migration project, we can assume that this graph is connected. We now use the standard Erdős-Rényi model [5], where this graph is constructed by connecting the components independently at random with a fixed probability $p$. One can ensure that the result is acyclic by defining any linear order on the vertices and include a directed edge only if it obeys the selected ordering. Each such edge is included in the graph with probability $p_{n}$, independently from every other edge. For such graphs, we have by Theorem 4.1 in [10], a variant of Theorem 1 in [4] that for such graphs, for growing $n$, the probability that $G$ is connected is given by
    $\begin{cases}0&\text{, if $c_{n}\to-\infty$}\\ e^{-e^{-c}}&\text{, if $c_{n}\to c$ constant}\\ 1&\text{, if $c_{n}\to\infty$}\\ \end{cases}$ (1)
    where $p_{n}=\frac{\mathop{\text{log}}\nolimits n+c_{n}}{n}$. Note that we can directly apply this Theorem to our directed acyclic graph situation, since such a graph is connected if and only if the undirected variant of it is connected.
    As our goal is to prove a lower bound on the number of edges in the condensation graph (with $n^{\prime}$ vertices!) in our model. We assume now the sparsest case $c_{n^{\prime}}\to c$ constant for which the graph is connected with high probability. Note that the doubly-exponential term in Equation (1) approaches $1$ already for comparatively small $c$ very quickly. For example, already for $c=10$, the value of $e^{-e^{-c}}\geq 0.99995$. But also for any $c>0$, we obtain for the number of edges in the condensation graph expectedly asymptotically
    $\frac{\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}}{n^{\prime}}\cdot\frac{n^{\prime}\cdot(n^{\prime}-1)}{2}\sim\frac{1}{2}n^{\prime}\mathop{\text{log}}\nolimits n^{\prime}.$
    That means that even in the sparsest case, we expect a considerable number of dependencies between the migration clusters. Furthermore, if we bound the number of components with a common ancestor of distance $k$, i. e. #$\#(v\to^{k}V)$ by a small function $f(n)=o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$, we obtain by Theorem 4 in Appendix -D for the condensation graph expectedly $d(G)\geq\frac{n}{f(n)\mathop{\text{log}}\nolimits n}$, that is, large chains of dependencies of the migration clusters.
This completes the proof of our main Theorem, proving that in a suitable sense we expect for any large migration project an expected intrinsic complexity with many small interdependent clusters. ∎
We thus showed that informally we have
Fact.
Any sufficiently large migration project has an intrinsic complexity due to many dependent (comparatively small) migration clusters. The migration of these clusters either takes many successive steps or includes at least one particularly difficult one.
[…]
TABLE I: Numerical evaluation of the expected behavior of migration clusters for migration projects with $n$ components rounded to the nearest integer. We see that one does not expect large mutual dependencies but the expected number of independent clusters is vast and the chain-dependencies are also large when we bound the number of migration clusters with fixed canonical migration time by $f(n)=\mathop{\text{log}}\nolimits n$ or $f(n)=\sqrt{n}/\mathop{\text{log}}\nolimits n$, respectively.
We conclude that even though the model does not represent the exact structure of every migration project, it gives a first sound basis for understanding the structure of migration projects in general. We summarize these results with a numerical evaluation of the relevant estimates in Table I with notation from Appendix -A.
V Practical Implications
Ott et al. define the term cryptographic agility as “the ability to make a transition from one cryptographic algorithm or implementation to another” [18]. Adopting cryptographic agility in that way ensures the flexibility to transition between algorithms or implementations in response to evolving threats. Thus, in order to implement cryptographic agility we need to prepare cryptographic migration. In turn, for preparing migration, we need a step-by-step migration plan. This plan must allow careful re-evaluation of dependencies and system integrity at each stage. This kind of approach is in line with making security solutions sustainable by allowing to adapt to changing threats and technologies. We have shown that the discussed migration approach based on graph models allows step-by-step evaluation and reevaluation throughout the whole process.
Furthermore, we have shown – as stated in Theorem 1 – big migration projects are intrinsically complex. These projects require many successive steps and in-depth strategic planning. Migration graphs can help in understanding this complexity involved and in considering the time needed to migrate, but also specify the number of components and the various dependencies. As we have seen, these dependencies can exist between single components, between interdependent clusters of components, and between external factors such as regulatory requirements, legal issues, or third-party components. The complexity of migrations requires expert knowledge and sufficient dedicated resources, resulting in significant costs. Economic implications and the risks of both, migrating and not migrating to PQC, must be evaluated. It is also crucial for stakeholders, policy makers, and non-technical decision makers to grasp this complexity and to develop appropriate organizational measures to address it. The model we propose can serve as a starting point to identify the highest priority parts of the system, e. g., by identifying the parts which are at the beginning of a migration chain, or which have the highest impact. We expect that from there, an iterative approach works best and we intend to investigate this in future work.
Given this general structure of practical migration, we think that our model can significantly support and enhance existing migration strategies and tools, particularly in the standardization of migration projects through guidelines, rules, and best practices. These models can optimize the process, provide practical examples for guidance, and reduce cost and effort. Initial steps towards such standardization include frameworks like CARAF, a risk management approach for post-quantum migration [12]. CARAF incorporates a threat and a risk assessment stage where our migration graphs can prove valuable. During threat assessment, migration clusters identified using our approach can be used to identify vulnerable subsystems. If one of the systems in the cluster is vulnerable, so is the cluster in its entirety. Likewise, during risk assessment, information about which systems need to be migrated together is helpful to establish cost and effort estimates. Consequently, we believe our approach complements and strengthens existing methods in the field, such as those explored by Näther et al. [14].
Currently, our approach is not supported by tools. We believe that the development and refinement of open-source tools and frameworks, along with ensuring easy access to these resources, can significantly lower the barriers for migration, especially for smaller organizations. These tools can streamline migration processes and offer cost-effective solutions for managing complex projects. We envision tooling to support the creation and maintenance of migration graphs as well as their analysis. Similar tools can be found in related areas such as network analysis and visualization, SBOMs and CBOMs, and graph-based databases. The formalization approach may facilitate the use and adaption of such tools by providing a structured understanding of the underlying complexities. Our future work will also address costs and effort associated with migrations and how these can be captured in the graphs.
In summary, managing the intrinsic complexity of large migration projects involves understanding the dependencies and strategically planning of the different migration steps, while keeping the economic implications in mind. Supporting companies with standardized guidelines and open-source tools is crucial, as is fostering an agile approach to sustainable security solutions. By addressing these key areas, organizations can overcome migration challenges and maintain robust security within their IT-infrastructures.
VI Conclusion
In this work, we used classical mathematical results from graph theory, probability theory, and combinatorial analysis to assess the difficulty of a very practical problem: Cryptographic migration of large IT-infrastructures. We formulated a formal model of cryptographic migratability in terms of directed graphs and showed several expected properties of these graphs. This approach not only elevates the discussion to a scientific level but also establishes a foundation for further research on more efficient migration strategies. In summary, we showed that migration projects always have an intrinsic complexity due to many dependent (comparatively small) migration clusters which (successively) depend on each other.
Still, there is a lot of work to be done, when it comes to practical migration projects: It is not clear how to obtain useful migration graphs in practice, such that they are neither trivial nor too vast. Analyzing how to model real-world IT-infrastructures or relevant subsets will clearly be a challenging task. Also, there is clearly room for improvements: As migration problems typically come with some kind of measurable cost per migration step, a refinement of our model would result in migration graphs where optimal migration strategies are not as easy to obtain as here. One further important open question is also the development of migration graphs and the changes in such graphs over time. Our work sets the stage for future research aimed at refining these models, integrating them with real-world data, and ultimately creating more adaptive, user-friendly tools that support real-world migration. We therefore aim here to scientifically substantiate migration approaches. By emphasizing the relevance of this formal approach, we hope to support the development of more robust migration strategies and (open-source) tooling that can keep pace with the evolving landscape of cryptographic standards. Future work should apply our model to real-world migration projects. The hope is to get a better insight of the complexity such projects not just in general, as discussed here, but for specific practically occurring cases. This might help to better understand real-world cryptographic migration processes, specifically in the context of PQC migration.
Acknowledgments
The authors thank Tobias J. Bauer for his valuable comments during the genesis of this article.
-A Migration Graphs
[…]
$v_{10}$$v_{11}$$v_{12}$$v_{9}$$v_{8}$$v_{7}$$v_{6}$$v_{5}$$v_{4}$$v_{2}$$v_{3}$$v_{1}$
Figure 2: An example of a simple migration graph.
-B Migration Clusters and Migratable Components
Definition 1 (Migration Cluster).
Let $G=(V,E)$ be a migration graph and $v\in V$. The migration cluster $c(v)$ of $v$ is set of all components $w\in V$ such that $w\in\mathop{\text{dep}}\nolimits(v)$ and $v\in\mathop{\text{dep}}\nolimits(w)$.
In the literature, the migration clusters are classically called the strongly connected components of $G$. For sake of exposition we stick to the term migration cluster in the context of cryptographic migration. Intuitively, the elements of $c(v)$ are precisely those components that have to be migrated at the same time as $v$ due to mutual dependencies. The migration clusters specify how complex single migration steps of a migration project will be. Additionally, the properties of the migration clusters inside a migration graph $G$ will shed light on the complexity of the resulting migration problem which we will describe below. In fact, we have the following important characteristics:
  1. 1.
    The size $s$ of the cluster $s(v):=\#c(v)$.
  2. 2.
    The degree $\mathop{\text{cdeg}}\nolimits(v):=\mathop{\text{argmax}}\nolimits_{w\in G|_{c(v)}}\mathop{\text{deg}}\nolimits(w)$.
Clearly, in the case $\mathop{\text{cdeg}}\nolimits(v)=0$ we have $s(v)=1$ since otherwise there would be two components $v_{1}\neq v_{2}\in c(v)$ with $v_{1}\in\mathop{\text{dep}}\nolimits(v_{2})$ and $v_{2}\in\mathop{\text{dep}}\nolimits(v_{1})$. But then $\mathop{\text{dep}}\nolimits(v_{1})=\mathop{\text{target}}\nolimits(v_{1}\to^{*}c(v_{2}))$ would be non-empty and thus $v_{1}\to^{*}c(v_{2})$ would also be non-empty. Therefore, we have $\mathop{\text{deg}}\nolimits_{G|_{c(v_{1})}}(v_{2})>0$, contradicting our assumption. Conversely, if $s(v)=1$ then $\mathop{\text{cdeg}}\nolimits(v)=0$ since $G|_{c(v)}$ consists of a single component only and has no edges. In practice this case occurs in the situation where we have a single component that simply can be migrated without any dependencies on other components. This is the easiest case in a migration process.
In the case $\mathop{\text{cdeg}}\nolimits(v)=1$ the migration cluster $c(v)$ is a cycle. In practice this case often occurs for $s(v)=2$ when there are two components that communicate using the same protocol. A migration of the protocol can only be done successfully, if both components are migrated at the same time. Cycles of larger sizes are uncommon in practice, cf. Section III-B. The set $m(v)$ of migratable components around $v$ is given by
$m(v)=\begin{cases}c(v)&\text{, if $c(v)\to\bar{c}(v)=\emptyset$}\\ \emptyset&\text{, otherwise}\\ \end{cases}$
where we write $\bar{c}(v)=V\setminus c(v)$. The set $m(v)$ specifies exactly which components can be migrated in $G$ when we migrate $v\in V$. If $v$ has dependencies outside its migration cluster, i. e. $v$ cannot be migrated, then $m(v)=\emptyset$. If there are no direct dependencies of $v$, then $m(v)=\left\{v\right\}$. Also, in all other cases $m(v)=c(v)$, i. e. the whole migration cluster. If $m(v)\neq\emptyset$, we call $v$ migratable (in $G$).
Now, in a migration project, there might be several subsequent steps to fully migrate all involved components, i. e. solve the migration problem the migration graph poses. We write $G^{(v)}=(V\setminus m(v),E\setminus(V\to m(v)))$ for the graph after a migration attempt of $v$. If $v$ is not migratable, then $G^{(v)}=G$. Generally, $G^{(v)}$ is exactly the graph restricted to the components outside the migratable components around $v$. Using similar to above the notation $\bar{m}(v)=V\setminus m(v)$, we can thus alternatively write $G^{(v)}=G|_{\bar{m}(v)}=(\bar{m}(v),\bar{m}(v)\to\bar{m}(v))$. We now show
Lemma 1.
In any (finite and non-empty) migration graph $G=(V,E)$ there is at least one component $v\in V$ that is migratable.
Proof.
Suppose there is no migratable component, i. e. $m(v)=\emptyset$ for all $v\in V$. By definition, $c(v)\to\bar{c}(v)$ is always non-empty, i. e. there is always a dependency to a component outside the migration cluster of $v$. Call such a dependency $w\in\bar{c}(v)$. Since $w$ is outside the migration cluster of $v$, $w$ cannot depend on elements of $c(v)$ and all dependencies of $w$ reside in the (strict) sub-graph $G|_{\bar{c}(v)}$. Repeating this process, after a finite number of steps, we end-up in a graph which only contains a single migration cluster $c(u)$ and $u\in V$ is migratable in $G$ contradicting our assumption. ∎
-C Migration Strategies
We will now define migration strategies of $G$. Formally, a strategy $S=(v_{1},\ldots,v_{\ell})\in V^{\ell}$ for $\ell\in\mathbb{N}$ is a sequence of components that are migrated one by one (of course together with their respective migration clusters). $\ell$ is the length of the strategy. For $1\leq i\leq\ell$ denote by $G_{i}$ the graph after successively migrating $v_{1},\ldots,v_{i}$, i. e. $G_{i}:=G_{i-1}^{(v_{i})}$, where $G_{0}=G$. The strategy $S$ is reasonable if for all $1\leq i\leq\ell$, we have $m(v_{i})\neq\emptyset$ in the graph $G_{i-1}$. That means that the component $v_{i}$ is in fact migratable in the graph $G_{i-1}$ and, therefore $G_{i}\subsetneq G_{i-1}$, i. e. the migration of $v_{i}$ yields a smaller graph. $S$ is successful, if $G_{\ell}$ is the empty graph. By Lemma 1 there is a reasonable, successful strategy for any migration graph $G$.
Lemma 2.
All reasonable, successful strategies $S=(v_{1},\ldots,v_{\ell})$ for a migration graph $G=(V,E)$ have the same length $\ell$.
Proof.
Consider two reasonable, successful strategies $S_{1}=(v_{1}^{1},\ldots,v_{\ell_{1}}^{1})$ and $S_{2}=(v_{1}^{2},\ldots,v_{\ell_{2}}^{2})$ with $\ell_{2}\geq\ell_{1}$. Since $S_{1}$ is successful, $V=\biguplus_{1\leq i\leq\ell_{1}}m(v_{i})$ partitions into the migration clusters given by $S_{1}$. Note that all $m(v_{i})=c(v_{i})$ are non-empty since $S_{1}$ is reasonable. Now, consider the (sub-)strategy $S_{3}=(v_{1}^{2},\ldots,v_{\ell_{1}}^{2})$ of the strategy $S_{2}$. By the partition property, there is for each $1\leq j\leq\ell_{1}$ an $1\leq i\leq\ell_{1}$ such that $v_{j}^{2}\in c(v_{i}^{1})$. Furthermore, for each $1\leq j_{1}<j_{2}\leq\ell_{1}$ and corresponding $1\leq i_{1},i_{2}\leq\ell_{1}$, we have $i_{1}\neq i_{2}$ since otherwise $v_{j_{2}}^{2}\in m(v_{j_{1}}^{2})\neq\emptyset$. The non-emptiness of $m(v_{j_{1}}^{2})$ follows from the assumption that $S_{2}$ is reasonable. But then $\sum_{1\leq j\leq\ell_{1}}s(v_{j}^{2})=\#V$ and the strategy $S_{3}$ is already successful. ∎
Due to Lemma 2 the migration length $\ell(G)$ as a property of the migration graph $G$ is well-defined. It basically expresses how many independent migration steps one has to perform in order to fully migrate a system. Equivalently, we see that the set of components $V$ partitions into the migration clusters of its components, i. e. $V=\bigcup_{v\in V}c(v)$ and for all $v_{1},v_{2}\in V$, we either have $c(v_{1})=c(v_{2})$ or $c(v_{1})\cap c(v_{2})=\emptyset$. In other words the relation $v_{1}\sim_{c}v_{2}:\Leftrightarrow c(v_{1})=c(v_{2})$ is an equivalence relation. Informally, we have
$c_{9}$$c_{8}$$c_{7}$$c_{6}$$c_{5}$$c_{4}$$c_{2}$$c_{3}$$c_{1}$
Figure 3: The condensation graph $G/c$ for the migration graph from Fig. 2.
[…]
Lemma 3.
For a migration graph $G=(V,E)$ the graph $G/c$ is acyclic.
Proof.
Assume $G/c$ is not acyclic, that is, there are $v_{1},v_{2}\in V$ such that $c(v_{1})\to^{+}c(v_{2})\in(E/c)^{+}$ and $c(v_{2})\to^{+}c(v_{1})\in(E/c)^{+}$ as usual denoting by the superscript $+$ the irreflexive transitive closure of the dependency relation. But then, by definition of $c$, we have $c(v_{1})=c(v_{2})$. ∎
[…]
Figure 4: The migration graph from Fig. 2 with its migration clusters and canonical migration times. In the example with $12$ components, we have $\ell(G)=9$ and $d(G)=5$.
[…]
  1. 1.
    $G$ partitions into the migration clusters of its components.
  2. 2.
    The migration clusters have only acyclic dependencies between each other.
  3. 3.
    The migration clusters have thus a canonical position in the migration process depending on how deep in the graph they are located.
Note that there are also non-canonical migration strategies, where a migration of a component $v\in V$ removed a dependency of another component $w\in V$ and even though there were other migratable components in the first place the migration of $w$ is done next.
-D Simple Trade-Offs
First of all, there is a simple trade-off between the number of migration clusters in a graph $G$ and the size of the clusters:
Theorem 2.
Assume that for a migration graph $G=(V,E)$ the migration length is bounded by a constant $k$, i. e. $\ell(G)\leq k$. Then, there is a component $v\in V$ with #$\#c(v)\geq n/k$.
Proof.
By definition $\ell(G)=\#(V/c)$ and $n=\sum_{c(u)\in V/c}\#c(u)$. Let $v$ be a component with a maximally large migration cluster, i. e. $v=\mathop{\text{argmax}}\nolimits_{u\in V}\#c(u)$. In general $v$ is not uniquely defined, but for our existential statement, this does not trouble us: Namely, for such $v\in V$, we have $n=\sum_{c(u)\in V/c}\#c(u)\leq\sum_{c(u)\in V/c}\#c(v)=\ell(G)\cdot\#c(v)\leq k\cdot\#c(v)$ and the claim follows. ∎
Also locally, there is a similar trade-off between the size of clusters and the number of independent components:
Theorem 3.
Consider a component $v$ in a migration graph $G$ with $m\in\mathbb{N}_{>0}$ direct dependencies. Assume there are at most $k$ independent clusters within those $m$ components. Then there is a migration cluster of size $s$ with $s\geq m/k$. Conversely, if each migration cluster has size at most $s$ then we have $k\geq m/s$ migration clusters.
Proof.
Assume each migration cluster has size $s<m/k$. Since there are at most $k$ independent clusters, we have at most $k\cdot s<m$ components. Contradiction. ∎
Thus, we see that we have to expect either many independent components or at least one large migration cluster. Another simple trade-off emerges, when we bound the number of descendants of any component in the graph suitably. Then, we get a large migration length as seen from
Theorem 4.
Let ${f\colon\mathbb{R}\rightarrow\mathbb{R}}$ be a smooth function in a real variable with $f\in o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$. Assume for any component $v$ in a migration graph $G=(V,E)$ with $n$ components, we have for each $k\geq 0$ the bound #$\#(v\to^{k}V)\leq f(n)$. Then $\ell(G)\geq n/f(n)$. In other words, the number of elements in the set of components with a common ancestor of distance $k$ has a trade-off with the length of the graph.
Proof.
Without loss of generality we can assume that $G$ has a single source $v$. Otherwise take the set of sources and consider the one with the longest path in $v\to^{*}V$. This path has length $\ell(G)$. For $0\leq k<\ell(G)$ define $W_{k}:=v\to^{k}V$. Since $G$ has only a single source, $\biguplus_{0\leq k<\ell(G)}W_{k}=V$, e. g. $n=\#V=\#\biguplus_{0\leq k<\ell(G)}W_{k}=\sum_{0\leq k<\ell(G)}\#W_{k}$. By assumption #$\#W_{k}\leq f(n)$ for all $0\leq k<\ell(G)$. Thus $n\leq\ell(G)\cdot f(n)$ and the claim follows. ∎
Note that the same trade-off exists for the condensed graph $G/c$ bounding the migration depth $d(G)$ of the graph $G$ from below when assuming for all $k$ bounds #$\#T(k)\leq f(n)$ on the number of migration clusters with canonical migration time $k$.